# KTaO3-Based Supercurrent Diode

### v1

@@ 3.1 Origin of the supercurrent diode effect in KTO WLs
The observed SDE in KTO WLs exhibits two characteristic behaviors: the effect reverses sign upon reversing the out-of-plane magnetic field and upon repositioning the WL from one edge to the other. These observations suggest that Meissner currents play a central role in the SDE mechanism. Previous studies of SDE in thin metallic superconducting films have highlighted the importance of Meissner currents and provide a framework for our analysis 21, 14. In two-dimensional superconducting systems, dissipation typically arises from vortex motion rather than Cooper pair breaking. To visualize vortex dynamics in our KTO WL geometry, we perform time-dependent Ginzburg-Landau (TDGL) simulations (see Methods). The simulated device consists of a vertical channel ($w=400\,\text{nm}$) with a narrow constriction near its left edge, constituting the WL (Figure 4(a)). Current source(drain) is defined at the top(bottom) edge of the channel. The simulated WL dimensions are $l_{\rm WL}=200\,\text{nm}$ and $w_{\rm WL}=50\,\text{nm}$, which are estimated based on Device F characteristics, as detailed in Supplementary Note S2. By design, it closely resembles Device C (Figure 1(c)) rotated counterclockwise by $90^{\circ}$.
@@ 3.2 Origin of the magnetic field enhancement of $I_{c}$ in Devices D and E
In Devices D and E, $I_{c}(B)$ exhibits two peaks (separated by $\Delta B$) where $I_{c}$ exceeds its zero-field value, roughly following a slanted M-shape. Previous studies reported enhancement of superconductivity by magnetic fields in Zn, MoGe, and Nb superconducting nanowires 29, 25. Specifically, Ref. 25 demonstrated an M-shaped $I_{c}(B)$ relation bearing great similarity to our observations, which was attributed to spin-exchange scattering between Cooper pairs and magnetic moments on the nanowire surfaces. We believe a similar mechanism can qualitatively explain our observations in KTO WLs. At $B=0$, local magnetic moments in the KTO sample induce exchange scattering of electrons, effectively breaking Cooper pairs and weakening superconductivity 23, 25, 24. Consequently, vortex entry and nucleation become more energetically favorable, suppressing $I_{c}$. At finite $B$, the magnetic moments are aligned by the field and exchange scattering is quenched, leading to higher $I_{c}$. At high $B$, conventional orbital and Zeeman effects dominate, weakening superconductivity and causing $I_{c}$ to decrease. Competition between magnetic moment pair breaking and orbital and Zeeman pair breaking can produce a non-monotonic, M-shaped $I_{c}(B)$ curve. Superimposed with SDE, this yields the slanted M-shaped $I_{c}(B)$ curves observed.
The fact that magnetic moments are distributed unevenly in the KTO sample explains the absence of M-shape in $I_{c}(B)$ of Devices A-C. Several reports 30, 31 have suggested magnetism in the LAO/KTO system, which is ascribed to Ta 5d states and oxygen vacancies. Magnetic impurities such as Fe and Ni can also be brought into the sample, especially during the polishing process of the crystal 32. Electrostatic gating with backgate ($V_{bg}$) is performed on Device D and E (Figure S13 and Figure S14). The magnitude of SDE $\eta_{max(min)}$ is barely affected by $V_{bg}$, always maxing out at $\approx 10\%$ (top panel, Figure S15), similar to the case of Devices A-C (discussed in Supplementary Note S1). Meanwhile, $\Delta B$ that separates the two $I_{c}$ peaks increases monotonically as $V_{bg}$ decreases (bottom panel, Figure S15), which suggests exchange scattering may be more pronounced with lower superfluid density and higher disorder. Enhancement of $I_{c}$ by field persists at temperature $T=500$ mK (Device D, Figure S16(b)). At $T=900$ mK however, $I_{c}(B)$ loses its M shape and only SDE can be seen (Figure S16(e)). Further increase in $T$ greatly suppresses superconductivity and smears out the SDE (Figure S16(h)). This temperature dependence indicates the onset of exchange scattering may occur at a temperature $\in(500\,\mathrm{mK},900\,\mathrm{mK})$, lower than the superconducting critical temperature $T_{c}$.
@@ 4 Outlook
The demonstration of a geometrically engineered, reconfigurable supercurrent diode effect (SDE) at the LAO/KTO interface establishes a uniquely versatile platform for both fundamental physics and quantum technologies. The ability to write, erase, and rewrite weak links with c-AFM lithography enables the rapid prototyping of non-reciprocal circuit elements for dissipationless electronics. This precise geometric control also enables on-demand engineering of energy landscape of vortices, making the system an ideal laboratory for systematic studies of 2D vortex dynamics. For example, a vortex pinning center may be defined by simply engaging a negatively-biased AFM tip on the device. Furthermore, large permittivity of KTO substrate enables electrostatic gating as a convenient tuning nob for devices. This, along with the large kinetic inductance of LAO/KTO interface 33, results in slow light speed in the system which is crucial for compact circuit elements. This work positions the LAO/KTO system at the forefront of research into 2D SDE, vortex physics, and the next generation of quantum circuits.
@@ 5 Methods
Time-dependent Ginzburgh Landau simulations
TDGL simulation is performed with the pyTDGL package (https://py-tdgl.readthedocs.io, Ref.34), which we modified to incorporate thermal fluctuations. In the pyTDGL package we first choose the following parameters for LAO/KTO(111) interface: Ginzburgh-Landau correlation length $\xi_{GL}=20\,\text{nm}$, normal state conductivity $\sigma=0.3\,\mathrm{S/\mu m}$, London penetration depth $\lambda_{L}=2.5\,\mathrm{\mu m}$, thickness of the LAO/KTO 2d electron gas (2DEG) $d=5$ nm and reduced temperature $t=T/T_{c}=0.05$. We justify these choices in the following:

### v2

@@ 3.1 Origin of the supercurrent diode effect in KTO WLs
The observed SDE in KTO WLs exhibits two characteristic behaviors: the effect reverses sign upon reversing the out-of-plane magnetic field and upon repositioning the WL from one edge to the other. These observations suggest that Meissner currents play a central role in the SDE mechanism. Previous studies of SDE in thin metallic superconducting films have highlighted the importance of Meissner currents and provide a framework for our analysis 21, 14. In two-dimensional superconducting systems, dissipation typically arises from vortex motion rather than Cooper pair breaking. To visualize vortex dynamics in our KTO WL geometry, we perform time-dependent Ginzburg-Landau (TDGL) simulations (see Methods and Ref.29). The simulated device consists of a vertical channel ($w=400\,\text{nm}$) with a narrow constriction near its left edge, constituting the WL (Figure 4(a)). Current source(drain) is defined at the top(bottom) edge of the channel. The simulated WL dimensions are $l_{\rm WL}=200\,\text{nm}$ and $w_{\rm WL}=50\,\text{nm}$, which are estimated based on Device F characteristics, as detailed in Supplementary Note S2. By design, it closely resembles Device C (Figure 1(c)) rotated counterclockwise by $90^{\circ}$.
@@ 3.2 Origin of the magnetic field enhancement of $I_{c}$ in Devices D and E
In Devices D and E, $I_{c}(B)$ exhibits two peaks (separated by $\Delta B$) where $I_{c}$ exceeds its zero-field value, roughly following a slanted M-shape. Previous studies reported enhancement of superconductivity by magnetic fields in Zn, MoGe, and Nb superconducting nanowires 30, 25. Specifically, Ref. 25 demonstrated an M-shaped $I_{c}(B)$ relation bearing great similarity to our observations, which was attributed to spin-exchange scattering between Cooper pairs and magnetic moments on the nanowire surfaces. We believe a similar mechanism can qualitatively explain our observations in KTO WLs. At $B=0$, local magnetic moments in the KTO sample induce exchange scattering of electrons, effectively breaking Cooper pairs and weakening superconductivity 23, 25, 24. Consequently, vortex entry and nucleation become more energetically favorable, suppressing $I_{c}$. At finite $B$, the magnetic moments are aligned by the field and exchange scattering is quenched, leading to higher $I_{c}$. At high $B$, conventional orbital and Zeeman effects dominate, weakening superconductivity and causing $I_{c}$ to decrease. Competition between magnetic moment pair breaking and orbital and Zeeman pair breaking can produce a non-monotonic, M-shaped $I_{c}(B)$ curve. Superimposed with SDE, this yields the slanted M-shaped $I_{c}(B)$ curves observed.
The fact that magnetic moments are distributed unevenly in the KTO sample explains the absence of M-shape in $I_{c}(B)$ of Devices A-C. Several reports 31, 32 have suggested magnetism in the LAO/KTO system, which is ascribed to Ta 5d states and oxygen vacancies. Magnetic impurities such as Fe and Ni can also be brought into the sample, especially during the polishing process of the crystal 33. Electrostatic gating with backgate ($V_{bg}$) is performed on Device D and E (Figure S13 and Figure S14). The magnitude of SDE $\eta_{max(min)}$ is barely affected by $V_{bg}$, always maxing out at $\approx 10\%$ (top panel, Figure S15), similar to the case of Devices A-C (discussed in Supplementary Note S1). Meanwhile, $\Delta B$ that separates the two $I_{c}$ peaks increases monotonically as $V_{bg}$ decreases (bottom panel, Figure S15), which suggests exchange scattering may be more pronounced with lower superfluid density and higher disorder. Enhancement of $I_{c}$ by field persists at temperature $T=500$ mK (Device D, Figure S16(b)). At $T=900$ mK however, $I_{c}(B)$ loses its M shape and only SDE can be seen (Figure S16(e)). Further increase in $T$ greatly suppresses superconductivity and smears out the SDE (Figure S16(h)). This temperature dependence indicates the onset of exchange scattering may occur at a temperature $\in(500\,\mathrm{mK},900\,\mathrm{mK})$, lower than the superconducting critical temperature $T_{c}$.
@@ 4 Outlook
The demonstration of a geometrically engineered, reconfigurable supercurrent diode effect (SDE) at the LAO/KTO interface establishes a uniquely versatile platform for both fundamental physics and quantum technologies. The ability to write, erase, and rewrite weak links with c-AFM lithography enables the rapid prototyping of non-reciprocal circuit elements for dissipationless electronics. This precise geometric control also enables on-demand engineering of energy landscape of vortices, making the system an ideal laboratory for systematic studies of 2D vortex dynamics. For example, a vortex pinning center may be defined by simply engaging a negatively-biased AFM tip on the device. Furthermore, large permittivity of KTO substrate enables electrostatic gating as a convenient tuning nob for devices. This, along with the large kinetic inductance of LAO/KTO interface 34, 35, results in slow light speed in the system which is crucial for compact circuit elements. This work positions the LAO/KTO system at the forefront of research into 2D SDE, vortex physics, and the next generation of quantum circuits.
@@ 5 Methods
Time-dependent Ginzburg Landau simulations
TDGL simulation is performed with the pyTDGL package (https://py-tdgl.readthedocs.io, Ref.29), which we modified to incorporate thermal fluctuations. In the pyTDGL package we first choose the following parameters for LAO/KTO(111) interface: Ginzburg-Landau correlation length $\xi_{GL}=20\,\text{nm}$, normal state conductivity $\sigma=0.3\,\mathrm{S/\mu m}$, London penetration depth $\lambda_{L}=2.5\,\mathrm{\mu m}$, thickness of the LAO/KTO 2d electron gas (2DEG) $d=5$ nm and reduced temperature $t=T/T_{c}=0.05$. We justify these choices in the following: